# Improved Bug Localization with AI Agents Leveraging Hypothesis and Dynamic Cognition

### v1

@@ Abstract.
Software bugs cost technology providers (e.g., AT & T) billions annually and cause developers to spend roughly 50% of their time on bug resolution. Traditional methods for bug localization often analyze the suspiciousness of code components (e.g., methods, documents) in isolation, overlooking their connections with other components in the codebase. Recent advances in Large Language Models (LLMs) and agentic AI techniques have shown strong potential for code understanding, but still lack causal reasoning during code exploration and struggle to manage growing context effectively, limiting their capability. In this paper, we present a novel agentic technique for bug localization –CogniGent– that overcomes the limitations above by leveraging multiple AI agents capable of causal reasoning, call-graph-based root cause analysis and context engineering. It emulates developers-inspired debugging practices (a.k.a., dynamic cognitive debugging) and conducts hypothesis testing to support bug localization. We evaluate CogniGent on a curated dataset of 591 bug reports using three widely adopted performance metrics and compare it against six established baselines from the literature. Experimental results show that our technique consistently outperformed existing traditional and LLM-based techniques, achieving MAP improvements of 23.33-38.57% at the document and method levels. Similar gains were observed in MRR, with increases of 25.14-53.74% at both granularity levels. Statistical significance tests also confirm the superiority of our technique. By addressing the reasoning, dependency, and context limitations, CogniGent advances the state of bug localization, bridging human-like cognition with agentic automation for improved performance.
@@
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: 34th IEEE/ACM International Conference on Program Comprehension (ICPC 2026), Research Track; 12–13 April 2026; Rio de Janeiro, Brazilisbn: 978-1-4503-XXXX-X/2026/04ccs: Software and its engineering Software maintenance toolsccs: Software and its engineering Maintaining softwareccs: Software and its engineering Software testing and debugging
@@ 1. Introduction
Traditional methods for bug localization can be grouped into two main categories: spectrum-based and Information Retrieval (IR)-based approaches. Spectrum-based techniques rely on execution traces, which are not always available in practice (Samir and Rahman, 2025a; Wang et al., 2015b; Moreno et al., 2014). In contrast, IR-based approaches attempt to localize bugs by matching tokens in bug reports with those in the source code (Wang et al., 2015a; Nguyen et al., 2011; Le et al., 2015; Lam et al., 2017b). While IR-based methods are lightweight and fast, they rely on surface-level token matching, which makes them prone to vocabulary mismatch between natural language and source code (Furnas et al., 1987), especially due to the variability in bug reports. To address the limitations, several avenues have been explored, such as query reformulations (Rahman and Roy, 2021; Sisman and Kak, 2012; Rahman et al., 2021; Chaparro et al., 2017), incorporating statistical insights from code change history, and leveraging past bug fixes (Youm et al., 2015; Wang and Lo, 2014). However, these techniques have not delivered significant improvements over earlier methods (Lee et al., 2018) and do not capture the deeper semantic context (e.g., program semantics (Ding et al., 2024)) needed for accurate bug localization (Samir and Rahman, 2025a; Qi et al., 2022).
@@
We evaluate our technique using 591 bug reports curated from an existing dataset of Samir et al. (Samir and Rahman, 2025b) and determine its performance at both the method and document levels using three widely adopted metrics: Mean Average Precision (MAP), Mean Reciprocal Rank (MRR), and HIT@K. We compared our approach with six baseline techniques (Foundation, 2021; Rahman and Roy, 2018; Saha et al., 2013; Samir and Rahman, 2025a; Xia et al., 2024; Chen et al., 2025) from two major areas—IR and LLM. CogniGent consistently outperformed existing traditional and LLM-based techniques, achieving MAP improvements of 23.33-38.57% at the document and method levels. Similar gains were observed in MRR, with increases of 25.14-53.74% across the same levels. These results underscore the effectiveness and superiority of CogniGent in software bug localization.
@@
  • A replication package (Anonymous, ) containing the prototype, curated dataset, and configuration details for third-party use.
@@
Title: The failsafe snapshot should be deleted after rollback successfully. Description: When a table exists and is in a disabled state, HBase supports restoring from a snapshot. In this scenario, HBase creates a failsafe snapshot for the disabled table and rolls back using it when restore fails. However, the failsafe snapshot remains even after successful rollback, and it should be deleted.
Technique Approach Rank Baseline IR (Lucene (Foundation, 2021)) Textual matching (Bug Report) 87 Agentless (Xia et al., 2024) Project structure cue + Semantic similarity 35 LocAgent (Chen et al., 2025) Agentic 8 CogniGent Dynamic Cognitive Debugging 1
@@ 2. Motivating Example
If we examine the source code in Fig. 1, at first glance, it appears correct—it calls deleteSnapshot() method after restoration, suggesting proper cleanup operation. However, given the reported bug, a developer might assume that the deletion may not be taking place after rollback. To verify this, they might navigate (e.g., Ctrl+Click) to the deleteSnapshot() method (line 20) for a close inspection, only to find it syntactically and logically correct. Given the role of the rollback operation, the developer might then inspect internalRestoreSnapshotAsync() method (line 3) to check whether it handles cleanup internally. As shown in Fig. 2, the method only initiates the restore or rollback operation but contains no deletion logic. The code contains two rollback operations within two-level nested try-catch blocks, complicating the control flows. When the first restore operation fails, an exception is caught, and a second attempt at restoration is made. When the second attempt to restoration is successful, a new RestoreSnapshotException is rethrown unexpedly (line 13). This rethrow causes control flow to exit the method before reaching the deleteSnapshot() (line 20) placed after the try–catch block. Even if takeFailSafeSnapshot is set to true, the deletion will not take place, and the failsafe snapshot will remain intact.
@@ 3.1. Design of AI Agents
In our agentic workflow, six role-specific agents form a pipeline to support bug localization: a restructuring agent (reorganize a bug report), a retrieval agent (collect candidate code segments), a filtering agent (filter out irrelevant candidates), a hypothesis agent (generate hypotheses), an investigation agent (test hypotheses), and an observer agent (validate hypotheses). We implement this workflow using the LangGraph library (LangChain, ), combining varying sizes of LLMs (e.g., Devstral (AI and AI, )) to balance efficiency, cost, and performance. The agents communicate through a shared LangGraph state (LangChain, Inc., ), where each agent uses only the parts relevant to its role and updates only the portion it contributes (e.g., adding hypotheses).
Since prompt design is central to the effectiveness of the workflow, we construct each agent’s system prompt following established prompt-engineering practices (White et al., 2023) and employ few-shot chain-of-thought prompting (Liang et al., 2023; Wei et al., 2022). We also apply a meta-prompting approach (Zhou et al., 2022) and iteratively refine the prompts with the help of LLMs (e.g., ChatGPT).
@@ 3.3.3. Filtering Documents
retrieveCandidates Retrieval Agent Retrieve the potential buggy source code segments (e.g., methods) from the bug reports.
exploreCallChain Supervisor Agent Determines and retrun the most relevant chain of code segment(s).
@@ 3.4. Dynamic Cognitive Debugging
In our proposed technique, we apply Dynamic Cognitive Debugging (DCD) to bug localization. Prior studies identify debugging as a cognitively demanding task that requires programmers to construct and refine mental models to locate and fix errors (Xu and Rajlich, 2004; Yoon and Garcia, 1998; Hu et al., 2024; Leite and Travassos, 2023). This process involves forming hypotheses, navigating the codebase to test them, and confirming or discarding those hypotheses based on relevant evidence. To emulate this important cognitive process, we leverage the reasoning capability of LLMs as AI agents, and employ them for code navigation and bug detection as follows.
@@ 3.4.1. Hypothesis Generation
During bug resolution, human experts typically form multiple hypotheses about the possible causes of a bug based on the bug report, available code and then investigate further (Xu and Rajlich, 2004). Similarly, we provide the bug report and the filtered code segments above (Step 6) to a powerful LLM (e.g., Devstral 20B parameters), and leverage the model to generate multiple competitive hypotheses (Step 7, Fig. 2). For each code segment, we guide the LLM to generate a hypothesis on the root cause of the reported bug, along with a confidence category (e.g., high, medium, or low) and a confidence score (e.g., 0-1). These categories represent the relative likelihood of each segment being buggy compared to the other candidates. We then retain the segments with high or medium confidence for deeper investigation in the subsequent steps.
@@ 3.4.2. Investigation: Hypothesis Testing using AI Agents
In this step, we simulate how developers investigate and navigate code during debugging (Step 8, Fig. 2). For example, when a bug report mentions a race condition, a developer may reason about it, inspect code related to threading functionality and then follow the relevant call chain for closer examination. Similarly, our technique applies contextual reasoning and autonomously determine whether the current segment provides sufficient evidence or whether parts of it (e.g., invoked methods) should be explored further.
@@ 14     $C^{*}\leftarrow(path,\mathcal{C}_{LLM}(seg))$
3.4.3. Observer
Once all investigations are complete, an independent observer agent evaluates each candidate explanation—either a single code segment or a call chain—along with its associated hypothesis and reasoning traces (Step 8, Fig. 2). The observer assesses how well each hypothesis is supported by the corresponding code and reasoning, and assigns a confidence score. The final suspiciousness score combines the supervisor’s reasoning confidence with the observer’s validation score, mitigating bias from either stage to find the most plausible root cause.
3.5. Final Ranking
4. Experiments
We curated a dataset of 591 bug reports from the work of Samir et al.(Samir and Rahman, 2025b), who extended the Bench4BL dataset (Lee et al., 2018) with recent reports. For evaluation, we employ three widely used metrics from the literature: Mean Average Precision (MAP), Mean Reciprocal Rank (MRR), and HIT@K (K=1, 5, 10). We experiment with three different LLMs and compare our solution, CogniGent, against five relevant baseline techniques to demonstrate its effectiveness. Through these experiments, we address three research questions:
  • RQ2 How do the individual agentic modules of CogniGent contribute to its overall performance in bug localization?
  • RQ3 Can CogniGent outperform the relevant baseline techniques in bug localization?
4.1. Dataset Construction
Recent advances in agent-based solutions have motivated the construction of new datasets (e.g.,  (Jimenez et al., 2023)). However, these datasets were primarily created for issue-fixing tasks and contain not only bug reports but also feature requests, questions, and other tasks (Herzig et al., 2013). Such a mixture of issues can negatively affect the performance of bug localization techniques (Herzig et al., 2013). Therefore, we curated a dataset composed exclusively of bug reports for our investigation.
@@ 4.2. Evaluation Metrics
Precision@K measures the accuracy of retrieved results up to rank-K for each occurrence of a buggy source code component (e.g., method or document) in the ranked list. Average Precision is the mean of these Precision@K values across all relevant (buggy) items for a given query. Consequently, the Mean Average Precision (MAP) is obtained by averaging the Average Precision scores across all queries Q in the dataset.
@@ 4.4. Evaluating CogniGent
Existing work has categorized bug reports into three categories based on the presence of structured entities in them – Stack Trace (ST), Program Elements (PE), and Natural Language (NL) types (Rahman and Roy, 2018). We categorize the reports using regular expressions (regex) (Rahman and Roy, 2018). To demonstrate the robustness of our technique, we also evaluated its performance across these bug types. Fig. 3 shows CogniGent’s bug localization performance in terms of MAP and MRR. Our technique exhibits consistent trends across all variants, performing best on PE and NL reports while struggling with ST reports at the document level. For ST-type reports, CogniGent (Devstral) achieves a MAP of 0.31—about 34% higher than baseline IR and up to 14.28% higher than the other two variants. A similar pattern is observed in MRR, confirming that bug reports with stack traces remain the most difficult to localize. In contrast, the Devstral and Qwen-Coder variants report higher gains for PE and NL, performing comparably with MAP and MRR improvements of 17.15% and 19.09%, respectively. A similar trend appears at the method level, though overall scores decline due to finer granularity. CogniGent (Devstral) remains dominant across all types of bug reports, achieving MAP values of 0.28 for PE and 0.22 for NL, while the Qwen-Coder variant trails by 4–6%. CogniGent (LLaMA) shows a 25–40% drop compared to the other two variants but still improves over baseline IR by up to 0.34% for ST and 0.32% for NL, with a slight decrease for PE. Thus, despite the increased difficulty of method-level localization, CogniGent presents a better alternative than the baseline for bug localization.
@@
Figure 4. Performance Across Single vs. Multi-Component Bugs
Given the prevalence of propagated bugs, we examined CogniGent’s performance on bugs spanning multiple source components (e.g., methods, documents). We first grouped bug reports by their ground-truth counts—1, 2, 3, 4, or 5+ buggy documents or segments (e.g., methods) —and evaluated our techniques’ performance using Mean Average Precision (MAP) at both document and method levels. Fig. 4 shows CogniGent’s performance across these groups of bug reports. CogniGent (Devstral) achieves a MAP of 0.40 for single-document bugs, improving to 0.43 and 0.46 for two and three documents (6.45% and 9.42% gains), but drops to 0.37 and 0.34 for four and five or more. Across all cases, it outperforms the IR baseline by 7.36–58.99%. Qwen-Coder follows a similar trend—starting at 0.39, surpassing Devstral by 4.6% for two-document bugs—and reaches up to 59.51% improvement over IR. LLaMA variant reports lower MAP scores overall but still achieves 0.28–19% gains over IR. At the method level, scores decline as bugs span more segments. CogniGent (Devstral) remains best, starting at 0.28 for single-segment bugs and dropping to 0.14–0.15 for broader scopes, with up to 39.70% improvement over IR. Qwen-Coder trails by 2–4%, while LLaMA drops to 0.12 for large bugs yet maintains 1.87–6.38% gains in other cases. Overall, while bug spread increases the difficulty of method-level reasoning, our technique effectively handles such cases through inter-component exploration and reasoning, achieving consistently strong localization performance.
Answering RQ2 - Contribution of Different Modules
Table 5. Effect of Individual Components in CogniGent
@@
Answering RQ3 - Comparison with Baseline Technique
We also compare CogniGent against three LLM-based techniques- BRaIn (Samir and Rahman, 2025a), Agentless (Xia et al., 2024) and LocAgent (Chen et al., 2025). BRaIn is a state of the art technique that improves bug localization through Intelligent Relevance Feedback. It captures contextual feedback from LLMs on initially retrieved documents, leverages them to enhance the query, and then employs Information Retrieval to localize bugs. For experiment, we collected the authors’ replication package (Asif Samir, 2024). AgentLess is a recent technique that combines LLM-based reasoning with embedding-based retrieval to identify buggy source documents, which are then used for fixing corresponding issues. It builds a structured repository representation, retrieves semantically related documents using embeddings, and then uses the LLM to reason over their skeleton forms functions to refine and rank the most likely buggy locations. We adapt the replication package from the authors (OpenAutoCoder, 2024) and use the localization portion of the technique for the experiment. On the other hand, LocAgent is a graph-guided agentic framework for issue localization. It builds a heterogeneous code graph of a project and creates text indexes (e.g., BM25) over code entities to quickly retrieve those relevant to a bug report. The LLM agent then searches and traverses the graph, reasoning over the retrieved entities to produce a ranked list of suspicious documents and functions. We collected the replication packages from the authors (Gerstein Lab, ) and adapt it for our dataset and version based localization and comparison.
Table 6 summarizes the comparison results between CogniGent and six baseline methods. Among traditional techniques, BLUiR achieves a MAP of 0.344, while BLIZZARD and the baseline-IR reach 0.379 and 0.330, respectively. CogniGent outperforms all of them with a MAP of 0.407, marking a up to 23.33% improvement, and achieves a maximum 25.14% gain in MRR (0.418). It also improves HIT@1 and HIT@5 to 0.374 and 0.475, yielding 16.31–47.33% and 5.00–7.25% gains against IR baselines in these metrics, respectively. At the method level, CogniGent achieves a MAP of 0.226 compared to 0.163, 0.191, and 0.177 for Baseline-IR, BLUiR, and BLIZZARD, showing 18.32–38.57% improvements. In MRR, HIT@1, and HIT@5, it also performs well, with gains of up to 53.74%, 86.11%, and 25.78%. These results demonstrate our technique’s benefits over the traditional baseline techniques.
@@
We further examine CogniGent’s performance on multi-component bugs, where a single bug affects multiple documents or methods. Fig. 5 compares CogniGent with BLIZZARD and LocAgent, the closest compettitors. In our dataset, we have 185 bug reports each of which triggers changes across multiple documents, totaling 523 buggy documents. Among these CogniGent localized 147 (28.10%) in the top-5 positions, outperforming LocAgent (132) and BLIZZARD (107). All three techniques commonly localized 54 documents, while CogniGent uniquely localized 57, compared to 37 by LocAgent and 24 by BLIZZARD. At the method level, 328 bug reports target 1,766 buggy methods in our dataset, where CogniGent localized 117 methods in the top-5, again outperforming LocAgent (102) and BLIZZARD (82). These results confirm that even at finer granularity, CogniGent provides consistently superior coverage of multi-component bugs.

### v2

@@ Abstract.
Software bugs cost technology providers (e.g., AT&T) billions annually and cause developers to spend roughly 50% of their time on bug resolution. Traditional methods for bug localization often analyze the suspiciousness of code components (e.g., methods, documents) in isolation, overlooking their connections with other components in the codebase. Recent advances in Large Language Models (LLMs) and agentic AI techniques have shown strong potential for code understanding, but still lack causal reasoning during code exploration and struggle to manage growing context effectively, limiting their capability. In this paper, we present a novel agentic technique for bug localization –CogniGent– that overcomes the limitations above by leveraging multiple AI agents capable of causal reasoning, call-graph-based root cause analysis and context engineering. It emulates developer-inspired debugging practices (a.k.a., dynamic cognitive debugging) and conducts hypothesis testing to support bug localization. We evaluate CogniGent on a curated dataset of 591 bug reports using three widely adopted performance metrics and compare it against six established baselines from the literature. Experimental results show that our technique consistently outperformed existing traditional and LLM-based techniques, achieving MAP improvements of 23.33-38.57% at the document and method levels. Similar gains were observed in MRR, with increases of 25.14-53.74% at both granularity levels. Statistical significance tests also confirm the superiority of our technique. By addressing the reasoning, dependency, and context limitations, CogniGent advances the state of bug localization, bridging human-like cognition with agentic automation for improved performance.
@@
journalyear: 2026copyright: acmlicensedconference: 34th IEEE/ACM International Conference on Program Comprehension; April 12–13, 2026; Rio de Janeiro, Brazilbooktitle: 34th IEEE/ACM International Conference on Program Comprehension (ICPC ’26), April 12–13, 2026, Rio de Janeiro, Brazildoi: 10.1145/3794763.3794818isbn: 979-8-4007-2482-4/2026/04ccs: Software and its engineering Software maintenance toolsccs: Software and its engineering Maintaining softwareccs: Software and its engineering Software testing and debugging
@@ 1. Introduction
Traditional methods for bug localization can be grouped into two main categories: spectrum-based and Information Retrieval (IR)-based approaches. Spectrum-based techniques rely on execution traces, which are not always available in practice (Samir and Rahman, 2025a; Wang et al., 2015b). In contrast, IR-based approaches attempt to localize bugs by matching tokens in bug reports with those in the source code (Wang et al., 2015a; Nguyen et al., 2011; Le et al., 2015; Lam et al., 2017b). While IR-based methods are lightweight and fast, they rely on surface-level token matching, which makes them prone to vocabulary mismatch between natural language and source code (Furnas et al., 1987), especially due to the variability in bug reports. To address the limitations, several avenues have been explored, such as query reformulations (Rahman and Roy, 2021; Sisman and Kak, 2012; Rahman et al., 2021; Chaparro et al., 2017), incorporating statistical insights from code change history, and leveraging past bug fixes (Youm et al., 2015; Wang and Lo, 2014). However, these techniques have not delivered significant improvements over earlier methods (Lee et al., 2018) and do not capture the deeper semantic context (e.g., program semantics (Ding et al., 2024)) needed for accurate localization of software bugs (Samir and Rahman, 2025a; Qi et al., 2022).
@@
We evaluate our technique using 591 bug reports curated from an existing dataset of Samir et al. (Samir and Rahman, 2025b) and determine its performance at both the method and document levels using three widely adopted metrics: Mean Average Precision (MAP), Mean Reciprocal Rank (MRR), and HIT@K. We compared our approach with six baseline techniques (Foundation, 2021; Rahman and Roy, 2018; Saha et al., 2013; Samir and Rahman, 2025a; Xia et al., 2024; Chen et al., 2025) from two major areas–IR and LLM. CogniGent consistently outperformed existing traditional and LLM-based techniques, achieving MAP improvements of 23.33-38.57% at the document and method levels. Similar gains were observed in MRR, with increases of 25.14-53.74% across the same levels. These results underscore the effectiveness and superiority of CogniGent in software bug localization.
@@
  • A replication package (15) containing the prototype, curated dataset, and configuration details for third-party use.
@@
Title: The failsafe snapshot should be deleted after rollback successfully.
Description: When a table exists and is in a disabled state, HBase supports restoring from a snapshot. In this scenario, HBase creates a failsafe snapshot for the disabled table and rolls back using it when restore fails. However, the failsafe snapshot remains even after successful rollback, and it should be deleted.
Technique Approach Rank
Baseline IR (Lucene (Foundation, 2021)) Textual matching (Bug Report) 87
Agentless (Xia et al., 2024) Project structure cue + Semantic similarity 35
LocAgent (Chen et al., 2025) Agentic 8
CogniGent Dynamic Cognitive Debugging 1
@@ 2. Motivating Example
If we examine the source code in Fig. 1, at first glance, it appears correct–it calls deleteSnapshot() method after restoration, suggesting proper cleanup operation. However, given the reported bug, a developer might assume that the deletion may not be taking place after rollback. To verify this, they might navigate (e.g., Ctrl+Click) to the deleteSnapshot() method (line 20) for a close inspection, only to find it syntactically and logically correct. Given the role of the rollback operation, the developer might then inspect internalRestoreSnapshotAsync() method (line 3) to check whether it handles cleanup internally. As shown in Fig. 2, the method only initiates the restore or rollback operation but contains no deletion logic. The code contains two rollback operations within two-level nested try-catch blocks, complicating the control flows. When the first restore operation fails, an exception is caught, and a second attempt at restoration is made. When the second attempt at restoration is successful, a new RestoreSnapshotException is rethrown unexpectedly (line 13). This rethrow causes control flow to exit the method before reaching the deleteSnapshot() (line 20) placed after the try–catch block. Even if takeFailSafeSnapshot is set to true, the deletion will not take place, and the failsafe snapshot will remain intact.
@@ 3.1. Design of AI Agents
In our agentic workflow, six role-specific agents form a pipeline to support bug localization: a restructuring agent (reorganizes a bug report), a retrieval agent (collects candidate code segments), a filtering agent (filters out irrelevant candidates), a hypothesis agent (generates hypotheses), an investigation agent (tests hypotheses), and an observer agent (validates hypotheses). We implement this workflow using the LangGraph library (LangChain, ), combining varying sizes of LLMs (e.g., Devstral (AI and AI, )) to balance efficiency, cost, and performance. The agents communicate through a shared LangGraph state (LangChain, Inc., ), where each agent uses and updates the relevant parts of the state (e.g., adding hypotheses).
Since prompt design is central to the effectiveness of our workflow, we construct an appropriate system prompt (Samir and Rahman, 2025a) for each agent following the best practices of prompt-engineering  (White et al., 2023; Samir and Rahman, 2025a) and employ few-shot chain-of-thought prompting (Liang et al., 2023; Wei et al., 2022). We apply meta-prompting (Zhou et al., 2022) to refine our prompts using LLMs (e.g., ChatGPT), as prompt quality can affect model performance (Wei et al., 2022; Zhou et al., 2022).
@@ 3.3.3. Filtering Documents
retrieveCandidates Retrieval Agent Retrieves the potential buggy source code segments (e.g., methods) from the bug reports.
exploreCallChain Supervisor Agent Determines and returns the most relevant chain of code segment(s).
@@ 3.4. Dynamic Cognitive Debugging
In our proposed technique, we apply Dynamic Cognitive Debugging (DCD) to bug localization. Prior studies identify debugging as a cognitively demanding task that requires programmers to construct and refine mental models to locate and fix errors (Xu and Rajlich, 2004; Yoon and Garcia, 1998; Hu et al., 2024; Leite and Travassos, 2023). This process involves forming hypotheses, navigating the codebase to test them, and confirming or discarding those hypotheses based on relevant evidence. To emulate this important cognitive process, we leverage the reasoning capability of LLMs as AI agents and employ them for code navigation and bug detection as follows.
@@ 3.4.1. Hypothesis Generation
During bug resolution, human experts typically form multiple hypotheses about the possible causes of a bug based on the bug report, available code, and then investigate further (Xu and Rajlich, 2004). Similarly, we provide the bug report and the filtered code segments above (Step 6) to a powerful LLM (e.g., Devstral 20B parameters), and leverage the model to generate multiple competitive hypotheses (Step 7, Fig. 2). For each code segment, we guide the LLM to generate a hypothesis on the root cause of the reported bug, along with a confidence category (e.g., high, medium, or low) and a confidence score (e.g., 0-1). These categories represent the relative likelihood of each segment being buggy compared to the other candidates. We then retain the segments with high or medium confidence for deeper investigation in the subsequent steps.
@@ 3.4.2. Investigation: Hypothesis Testing using AI Agents
In this step, we simulate how developers investigate and navigate code during debugging (Step 8, Fig. 2). For example, when a bug report mentions a race condition, a developer may reason about it, inspect code related to threading functionality and then follow the relevant call chain for closer examination. Similarly, our technique applies contextual reasoning and autonomously determines whether the current segment provides sufficient evidence or whether parts of it (e.g., invoked methods) should be explored further.
@@
3.4.3. Observer
Once all investigations are complete, an independent observer agent evaluates each candidate explanation–either a single code segment or a call chain—along with its associated hypothesis and reasoning traces (Step 8, Fig. 2). The observer assesses how well each hypothesis is supported by the corresponding code and reasoning, and assigns a confidence score. The final suspiciousness score combines the supervisor’s reasoning confidence with the observer’s validation score, mitigating bias from either stage to find the most plausible root cause.
3.5. Final Ranking
@@ 14     $C^{*}\leftarrow(path,\mathcal{C}_{LLM}(seg))$
4. Experiments
We curated a dataset of 591 bug reports from the work of Samir et al. (Samir and Rahman, 2025b), who extended the Bench4BL dataset (Lee et al., 2018) with recent reports. For evaluation, we employ three widely used metrics from the literature: Mean Average Precision (MAP), Mean Reciprocal Rank (MRR), and HIT@K (K=1, 5, 10). We experiment with three different LLMs and compare our solution, CogniGent, against six relevant baseline techniques to demonstrate its effectiveness. Through these experiments, we address three research questions:
  • RQ2: How do the individual agentic modules of CogniGent contribute to its overall performance in bug localization?
  • RQ3: Can CogniGent outperform the relevant baseline techniques in bug localization?
4.1. Dataset Construction
Recent advances in agent-based solutions have motivated the construction of new datasets (e.g., SWE-bench (Jimenez et al., 2023)). However, these datasets were primarily created for issue-fixing tasks and contain not only bug reports but also feature requests, questions, and other tasks (Herzig et al., 2013). Such a mixture of issues can negatively affect the performance of bug localization techniques (Herzig et al., 2013). Therefore, we curated a dataset composed exclusively of bug reports for our investigation.
@@ 4.2. Evaluation Metrics
Precision@K measures the accuracy of retrieved results up to rank-K for each occurrence of a buggy source code component (e.g., method or document) in the ranked list. Average Precision is the mean of these Precision@K values across all relevant (buggy) items for a given query. Consequently, the Mean Average Precision (MAP) is obtained by averaging the Average Precision scores across all queries. Q in the dataset.
@@ 4.4. Evaluating CogniGent
Existing work has categorized bug reports into three categories based on the presence of structured entities in them – Stack Trace (ST), Program Elements (PE), and Natural Language (NL) types (Rahman and Roy, 2018). We categorize the reports using regular expressions (regex) (Rahman and Roy, 2018). To demonstrate the robustness of our technique, we also evaluated its performance across these bug types. Fig. 3 shows CogniGent’s bug localization performance in terms of MAP and MRR. Our technique exhibits consistent trends across all variants, performing best on PE and NL reports while struggling with ST reports at the document level. For ST-type reports, CogniGent (Devstral) achieves a MAP of 0.31–about 34% higher than baseline IR and up to 14.28% higher than the other two variants. A similar pattern is observed in MRR, confirming that bug reports with stack traces remain the most difficult to localize. In contrast, the Devstral and Qwen-Coder variants report higher gains for PE and NL, performing comparably with MAP and MRR improvements of 17.15% and 19.09%, respectively. A similar trend appears at the method level, though overall scores decline due to finer granularity. CogniGent (Devstral) remains dominant across all types of bug reports, achieving MAP values of 0.28 for PE and 0.22 for NL, while the Qwen-Coder variant trails by 4–6%. CogniGent (LLaMA) shows a 25–40% drop compared to the other two variants but still improves over baseline IR by up to 0.34% for ST and 0.32% for NL, with a slight decrease for PE. Thus, despite the increased difficulty of method-level localization, CogniGent presents a better alternative than the baseline for bug localization.
@@
Figure 4. Performance for Single vs. Multi-Component Bugs
Given the prevalence of propagated bugs, we examined CogniGent’s performance on bugs spanning multiple source components (e.g., methods, documents). We first grouped bug reports by their ground-truth counts-1, 2, 3, 4, or 5+ buggy documents or segments (e.g., methods) –and evaluated our techniques’ performance using Mean Average Precision (MAP) at both document and method levels. Fig. 4 shows CogniGent’s performance across these groups of bug reports. CogniGent (Devstral) achieves a MAP of 0.40 for single-document bugs, improving to 0.43 and 0.46 for two and three documents (6.45% and 9.42% gains), but drops to 0.37 and 0.34 for four and five or more. Across all cases, it outperforms the IR baseline by 7.36–58.99%. Qwen-Coder follows a similar trend—starting at 0.39, surpassing Devstral by 4.6% for two-document bugs—and reaches up to 59.51% improvement over IR. LLaMA variant reports lower MAP scores overall but still achieves 0.28–19% gains over IR. At the method level, scores decline as bugs span more segments. CogniGent (Devstral) remains best, starting at 0.28 for single-segment bugs and dropping to 0.14–0.15 for broader scopes, with up to 39.70% improvement over IR. Qwen-Coder trails by 2–4%, while LLaMA drops to 0.12 for large bugs yet maintains 1.87–6.38% gains in other cases. Overall, while bug spread increases the difficulty of method-level reasoning, our technique effectively handles such cases through inter-component exploration and reasoning, achieving consistently strong localization performance.
Table 5. Effects of Individual Components in CogniGent
@@
Answering RQ2 - Contribution of Different Modules
Answering RQ3 - Comparison with Baseline Technique
@@
We also compare CogniGent against three LLM-based techniques- BRaIn (Samir and Rahman, 2025a), Agentless (Xia et al., 2024) and LocAgent (Chen et al., 2025). BRaIn is a state of the art technique that improves bug localization through Intelligent Relevance Feedback. It captures contextual feedback from LLMs on initially retrieved documents, leverages them to enhance the query, and then employs Information Retrieval to localize bugs. For the experiment, we collected the authors’ replication package (Asif Samir, 2024). AgentLess is a recent technique that combines LLM-based reasoning with embedding-based retrieval to identify buggy source documents, which are then used for fixing corresponding issues. It builds a structured repository representation, retrieves semantically related documents using embeddings, and then uses the LLM to reason over their skeleton forms functions to refine and rank the most likely buggy locations. We adapt the replication package from the authors (OpenAutoCoder, 2024) and use the localization portion of the technique for the experiment. On the other hand, LocAgent is a graph-guided agentic framework for issue localization. It builds a heterogeneous code graph of a project and creates text indexes (e.g., BM25) over code entities to quickly retrieve those relevant to a bug report. The LLM agent then searches and traverses the graph, reasoning over the retrieved entities to produce a ranked list of suspicious documents and functions. We collected the replication packages from the authors (Gerstein Lab, ) and adapted them for our dataset and version-based localization and comparison.
@@
Table 6 summarizes the comparison results between CogniGent and six baseline methods. Among traditional techniques, BLUiR achieves a MAP of 0.344, while BLIZZARD and the baseline-IR reach 0.379 and 0.330, respectively. CogniGent outperforms all of them with a MAP of 0.407, marking up to 23.33% improvement, and achieves a maximum 25.14% gain in MRR (0.418). It also improves HIT@1 and HIT@5 to 0.374 and 0.475, yielding 16.31–47.33% and 5.00–7.25% gains against IR baselines in these metrics, respectively. At the method level, CogniGent achieves a MAP of 0.226 compared to 0.163, 0.191, and 0.177 for Baseline-IR, BLUiR, and BLIZZARD, showing 18.32–38.57% improvements. In MRR, HIT@1, and HIT@5, it also performs well, with gains of up to 53.74%, 86.11%, and 25.78%. These results demonstrate our technique’s benefits over the traditional baseline techniques.
@@
We further examine CogniGent’s performance on multi-component bugs, where a single bug affects multiple documents or methods. Fig. 5 compares CogniGent with BLIZZARD and LocAgent, the closest competitors. In our dataset, we have 185 bug reports, each of which triggers changes across multiple documents, totaling 523 buggy documents. Among these CogniGent localized 147 (28.10%) in the top-5 positions, outperforming LocAgent (132) and BLIZZARD (107). All three techniques commonly localized 54 documents, while CogniGent uniquely localized 57, compared to 37 by LocAgent and 24 by BLIZZARD. At the method level, 328 bug reports target 1,766 buggy methods in our dataset, where CogniGent localized 117 methods in the top-5, again outperforming LocAgent (102) and BLIZZARD (82). These results confirm that even at finer granularity, CogniGent provides consistently superior coverage of multi-component bugs.